# Measuring Gravitational Wave Spectrum from Electroweak Phase Transition and Higgs Self-Couplings

### v2

@@ 4.3 Likelihood
In practical analysis, the amount of data may be too large to handle directly, which necessitates the use of approximation methods. Therefore, in the literature, two approximation methods are usually considered based on the central limit theorem (CLT) and coarse graining (CG) Caprini:2024hue. In this work, we use the full likelihood above directly.
5 Fisher Information Matrix and Bayesian Analysis
In this section, we investigate two complementary approaches to parameter estimation in GW data analysis: the FIM doi:https://doi.org/10.1002/9780470173862.ch4 and Bayesian inference Gregory_2005. The FIM provides an analytical framework for quantifying the precision of parameter estimation by evaluating the curvature of the likelihood function around its maximum. Under the assumption that the likelihood is well approximated by a multivariate Gaussian, the FIM yields theoretical lower bounds on parameter uncertainties, such as the Cramér–Rao lower bound (CRLB) doi:https://doi.org/10.1002/9780470173862.ch4. Owing to its computational efficiency, the FIM is particularly valuable for preliminary sensitivity studies and for visualizing parameter correlations via confidence ellipses.
In contrast, Bayesian inference, typically implemented through MCMC sampling gelman2013bayesian, provides a more general and robust characterization of the posterior distribution. Unlike the FIM, MCMC does not assume Gaussianity or linearity, making it especially suitable for exploring non-Gaussian, multimodal, or highly degenerate likelihood surfaces. By generating a representative set of samples from the posterior distribution, MCMC yields statistically accurate estimates of parameter uncertainties and correlations. In this study, we employ both analytical derivations and simplified toy models to compare these two methods, highlighting the regimes in which the FIM offers reliable approximations and the conditions under which full Bayesian inference becomes essential.
5.1 Fisher Information Matrix and the Cramér-Rao Lower Bound
The FIM serves as a fundamental method for quantifying the amount of information that an observed random variable contains about an unknown parameter. This concept plays a crucial role in evaluating the efficiency of estimators. In particular, efficient estimators—those that achieve the minimum possible variance—are generally preferred in statistical inference. More relevantly, the FIM is closely connected to the CRLB, which provides a theoretical lower limit on the variance of any unbiased estimator. This bound establishes a benchmark for the best achievable precision in parameter estimation, thereby offering a rigorous standard against which the performance of practical estimators can be assessed.
@@ 5.1 Fisher Information Matrix and the Cramér-Rao Lower Bound
In this section, we derive the general expression for the FIM. Before doing so, it is necessary to examine the analytical form of the log-posterior function, which forms the basis for the subsequent derivationGregory_2005:
$\ln\left[p(\bm{\theta})\mathcal{L}(\bm{\theta})\right]=\ln\left[p(\hat{\bm{\theta}})\mathcal{L}(\hat{\bm{\theta}})\right]-\frac{1}{2}\sum_{ij}\left(\theta_{i}-\hat{\theta}_{i}\right)(C^{-1})_{ij}\left(\theta_{j}-\hat{\theta}_{j}\right),$ (5.1)
@@
$F_{ij}\equiv(C^{-1})_{ij}=-E\left(\frac{\partial^{2}\ln p(\bm{\theta})\mathcal{L}(\bm{\theta})}{\partial\theta_{i}\partial\theta_{j}}\right),$ (5.2)
@@
$F_{ij}=F_{ij}^{\text{prior}}+F_{ij}^{\text{likelihood}},$ (5.3)
@@
$F_{ij}^{\text{prior}}=-E\left(\frac{\partial^{2}\ln p(\bm{\theta})}{\partial\theta_{i}\partial\theta_{j}}\right),\qquad F_{ij}^{\text{likelihood}}=-E\left(\frac{\partial^{2}\ln\mathcal{L}(\bm{\theta})}{\partial\theta_{i}\partial\theta_{j}}\right).$ (5.4)
For an unbiased estimator $\hat{\theta}_{i}$ of an unknown parameter, the CRLB imposes the following lower bound on its standard deviation doi:https://doi.org/10.1002/9780470173862.ch4:
$\Delta\hat{\theta}_{i}\geq\sqrt{(F^{-1})_{ii}}.$ (5.5)
@@
To visualize these correlations, one may construct confidence ellipses, which provide a geometric interpretation of the joint uncertainty between two parameters. These ellipses are derived from the second term in Eq. (5.1). The construction begins by analyzing the so-called Mahalanobis distance Mahalanobis:1936gd:
$d_{M}^{2}=\sum\limits_{ij}\left(\theta_{i}-\hat{\theta}_{i}\right)(C^{-1})_{ij}\left(\theta_{j}-\hat{\theta}_{j}\right),$ (5.6)
@@
$\Sigma=\begin{pmatrix}C_{\alpha\alpha}&C_{\alpha\beta}\\ C_{\alpha\beta}&C_{\beta\beta}\end{pmatrix}$, (5.7)
@@
$\begin{pmatrix}z_{\alpha}\\ z_{\beta}\end{pmatrix}=\Sigma^{-\frac{1}{2}}\begin{pmatrix}\theta_{\alpha}-\hat{\theta}_{\alpha}\\ \theta_{\beta}-\hat{\theta}_{\beta}\end{pmatrix}$, (5.8)
@@
5.2 Bayesian Inference and MCMC Sampling

### v3

@@ 4.3 Likelihood
In practical analysis, the amount of data may be too large to handle directly, which necessitates the use of approximation methods. Therefore, in the literature, two approximation methods are usually considered based on the central limit theorem (CLT) and coarse graining (CG) Caprini et al. (2024). In this work, we use the full likelihood above directly.
4.4 Statistical Inference Framework
To extract physical information from the simulated SGWB data, we employ two complementary statistical approaches: the FIM and Bayesian inference based on MCMC sampling.
The FIM provides a fast and computationally efficient method to estimate parameter uncertainties by approximating the likelihood function as a multivariate Gaussian in the vicinity of its maximum. Under this assumption, the inverse of the FIM yields the Cramér–Rao lower bound (CRLB), which represents the minimum achievable variance for any unbiased estimator. This method is particularly useful for forecasting sensitivities, estimating relative uncertainties, and visualizing parameter correlations through confidence ellipses.
In contrast, Bayesian inference does not rely on Gaussian or linear assumptions. Instead, it samples the full posterior probability distribution of the model parameters, allowing for a faithful characterization of non-Gaussian features, parameter degeneracies, and correlations. While computationally more demanding, MCMC sampling provides robust and statistically complete uncertainty estimates.
In this work, the FIM is primarily used for fast forecasts and qualitative understanding, while MCMC sampling is adopted to obtain the final and reliable parameter constraints. The detailed mathematical formulation and derivations underlying these statistical methods are presented in Appendix A.
5 Results
With the statistical analysis framework established in the preceding sections, we are now ready to investigate, via simulation, how the parameters in our model can be measured. This study is conducted at multiple levels of complexity. We begin with a simulation that includes only detector noise and a SGWB of astrophysical origin. Subsequently, we inject a SGWB from the SWs, as introduced in Sec. 3. The goal is to infer the model parameters and quantify the associated uncertainties. The inferred results are then mapped onto the xSM parameter space and used to extract physical observables, such as the Higgs self-couplings.
5.1 Measurements of the Astrophysical Parameters
where $\Omega_{\text{ast}}$ denotes the amplitude of the SGWB spectrum, and the reference frequency is chosen as $f_{\text{ref}}=1\,\text{mHz}$. The function $\mathcal{R}_{A,E}(f)$ represents the detector response function in the $A$ or $E$ channel. The corresponding likelihood function is given by Eq. (4.20).
Firstly, the FIM is employed to analyze the parameters in the spectrum, commonly denoted as $\bf\theta$} and to derive the corresponding confidence ellipses. By substituting Eq. (4.20) into Eq. (A.4), the FIM can be written in the following form Boileau et al. (2021a):
Figure 3: Comparison of parameter estimation uncertainties obtained from the FIM (blue) and MCMC sampling (red) for four representative parameters: two instrumental noise parameters, $N_{\mathrm{acc}}$ and $\delta x$, and two astrophysical parameters, $\Omega_{\mathrm{ast}}$ and $\varepsilon$. The injected values, $N_{\mathrm{acc}}=3\times 10^{-15}$, $\delta x=8\times 10^{-12}$, $\log_{10}\Omega_{\mathrm{ast}}=-11.5$, and $\varepsilon=2/3$, are indicated by solid blue lines, while the MCMC-recovered best-fit values are shown as solid red lines. The dark and light blue shaded regions denote the 68% and 95% confidence contours predicted by the FIM, respectively, whereas the red contours represent the corresponding credible regions from MCMC sampling. The diagonal panels display the marginalized one-dimensional posterior distributions from both approaches, with dashed vertical lines marking the $1\sigma$ intervals (blue for FIM and red for MCMC). The relative uncertainties predicted by the FIM are approximately 0.088%, 0.006%, 0.033%, and 0.738%, while those obtained from the MCMC posteriors are nearly identical: 0.088%, 0.006%, 0.033%, and 0.728%. The shape and orientation of the ellipses illustrate the correlations between parameters: elongated and tilted contours indicate strong degeneracies, whereas more circular contours suggest weaker coupling.
[…]
Fig. 3 presents a concrete example of such a comparison. We focus on four representative parameters in the SGWB model: two related to instrumental noises, $N_{\rm acc}$ and $\delta x$, and two associated with the astrophysical background, $\Omega_{\rm ast}$ and $\varepsilon$. The injected values are $N_{\rm acc}=3\times 10^{-15}$, $\delta x=8\times 10^{-12}$, $\log_{10}\Omega_{\rm ast}=-11.5$, and $\varepsilon=2/3$, indicated by the blue lines, while the red lines indicate the parameter values recovered from the MCMC analysis, corresponding to a SNR of 739. The 68% and 95% confidence regions predicted by the FIM are shown as dark and light blue ellipses, respectively, and the corresponding credible regions from MCMC sampling are displayed in red. In addition, the $1\sigma$ ranges from the marginalized one-dimensional posteriors are indicated by dashed lines, with blue for FIM and red for MCMC. From these plots, we find that the relative uncertainties estimated via the FIM, using the bound in Eq. (A.5), are approximately 0.088% for $N_{\rm acc}$, 0.006% for $\delta x$, 0.033% for $\log_{10}\Omega_{\rm ast}$, and 0.738% for $\varepsilon$. The corresponding uncertainties from the MCMC posteriors are nearly identical: 0.088%, 0.006%, 0.033%, and 0.728%. These results demonstrate high precision in parameter recovery across both approaches.
5.2 Measurements of the Thermodynamics Parameters
Here, in addition to the contributions from instrumental noises and the astrophysical SGWB, we incorporate the SGWB generated by the EWPT, specifically the dominant contribution from SWs discussed earlier. The likelihood function remains the same as in previous analyzes, with a modified signal model to include this additional component:
Using this extended signal model, we estimate the joint posterior distribution for six parameters: two instrumental noise parameters, $N_{\text{acc}}$ and $\delta x$; two astrophysical background parameters, $\Omega_{\text{ast}}$ and $\varepsilon$; and two parameters characterizing the phase transition of the SGWB, $\Omega_{0}$ and $f_{\mathrm{p}}$. The resulting confidence ellipses are shown in Fig. 4, providing a comprehensive view of the parameter correlations and uncertainties as derived from both the FIM and full posterior exploration via MCMC sampling within the Bayesian framework. This comparison allows us to assess the validity of the Gaussian approximation in a higher-dimensional, physically motivated parameter space. The injected values, used consistently in both simulation and inference, are $N_{\text{acc}}=3\times 10^{-15}$, $\delta x=8\times 10^{-12}$, $\log_{10}\Omega_{\rm ast}=-11.5$, $\varepsilon=2/3$, $\log_{10}\Omega_{0}=-12.0$, and $\log_{10}(f_{p}/\rm Hz)=-2.13$. With this set of parameters, the calculated SNRs for the astrophysical background and the SW contribution are 739 and 52, respectively. The relative uncertainties predicted by the FIM for these parameters are approximately 0.096%, 0.007%, 0.057%, 1.333%, 0.491%, and 1.29%, respectively. In contrast, the corresponding uncertainties estimated from MCMC sampling are very similar: 0.097%, 0.007%, 0.057%, 1.329%, 0.552%, and 1.406%.
At first glance, the MCMC-derived relative uncertainty for $\varepsilon$ appear slightly smaller than those predicted by the FIM, which may seem counterintuitive since the FIM is generally expected to provide the best forecasts under the assumption of a locally Gaussian posterior. This apparent tension can be understood by recalling that Fisher forecasts probe only the local curvature of the likelihood around the maximum-likelihood point. By construction, they neglect non-Gaussian features that are naturally captured in MCMC analyzes. In practice, the confidence ellipses from the FIM primarily illustrate the local shape of the posterior, whereas the standard deviations derived from MCMC chains account for the full posterior volume, including possible asymmetries, long tails, and mild non-linearities. As a result, the projected ellipses from MCMC can appear comparable or even slightly tighter, while the corresponding marginalized uncertainties remain broader. The inclusion of $\Omega_{0}$ and $f_{\mathrm{p}}$ adds further structure to the likelihood surface since these parameters directly determine the amplitude and peak frequency of the phase-transition signal. Their coupling with $\Omega_{\text{ast}}$ and $\varepsilon$ introduces moderate degeneracies, which broaden the confidence regions along certain directions in parameter space. Overall, while the FIM remains a valuable first-order approximation, its local Gaussian nature becomes less reliable as dimensionality and model complexity increase, underscoring the need for full Bayesian inference to obtain robust uncertainty quantification.
Figure 5: Relative uncertainty of $\log_{10}\Omega_{0}$, defined as $\Delta\log_{10}\Omega_{0}/\log_{10}\Omega_{0}$, as a function of $\log_{10}\Omega_{0}$ for different astrophysical background amplitudes $\log_{10}\Omega_{\text{ast}}$, derived using the FIM (lines) and MCMC for one scenario (points). The MCMC results (orange points) are consistent with the Fisher forecasts (orange line), both obtained for $\log_{10}\Omega_{\text{ast}}=-11.5$.
To further investigate how the detectability of the phase transition signal depends on the fiducial model parameters, we analyze the relative uncertainty of $\log_{10}\Omega_{0}$ as a function of its injected value. In particular, we also explore how this uncertainty is influenced by the presence of an overlapping astrophysical SGWB component, characterized by different values of $\log_{10}\Omega_{\text{ast}}$. Using the FIM formalism, we compute the relative error $\frac{\Delta\log_{10}\Omega_{0}}{\log_{10}\Omega_{0}}$ under three representative scenarios: $\log_{10}\Omega_{\text{ast}}=-10.5$, 11.5-11.5, and 12.5-12.5. The results, illustrated in Fig. 5, demonstrate that a stronger astrophysical background increases the uncertainty in $\Omega_{0}$, especially when $\Omega_{0}$ is small, as expected. This highlights the importance of disentangling different SGWB components when interpreting observational data. We have also shown in this figure the comparison between MCMC-derived uncertainties (with stars) and those from the FIM for one case, which demonstrates the overall agreement between the two approaches.
In the case of the full SWs model described by Eq. (3.1), parameter degeneracies naturally arise from the presence of three free thermodynamic parameters: $\beta/H_{\mathrm{n}}$, $T_{\mathrm{n}}$, and $K=\alpha/(1+\alpha)$. These degeneracies pose challenges for accurate parameter inference and interpretation. To mitigate this issue, we follow the strategy of Caprini et al. (2024), fixing $T_{\mathrm{n}}$ at random values drawn from a uniform distribution. The resulting confidence ellipses for all parameters are shown in Fig. 6, enabling a direct comparison between Fisher forecasts and posterior distributions. This approach provides insight into the attainable parameter constraints and shows the impact of degeneracies under the thermodynamic parameterization. Our results can be compared with the analysis of Caprini et al. (2024).
To summarize, the FIM provides a fast and computationally efficient way to forecast uncertainties under the assumption of a local Gaussian posterior. In contrast, MCMC sampling—though computationally more demanding—yields the full posterior distribution and faithfully captures non-Gaussian features and parameter degeneracies. In both cases, we observe overall agreement between the two approaches, and have thus obtained the measurement uncertainties for the parameters governing the GW spectrum and that of the phase transitions.
5.3 Measurements of xSM Model Parameters
Due to the aforementioned degeneracy, we adopt a strategy based on the directly inferred spectral parameters $(\log_{10}\Omega_{0},\log_{10}(f_{p}/\rm Hz))$ and identify the regions in the xSM parameter space that are consistent with these measurements. To illustrate how the resulting constraints propagate into the underlying particle-physics model, we project the scanned xSM parameter points onto the spectral parameter plane $(\log_{10}\Omega_{0},\log_{10}(f_{p}/\rm Hz))$. Fig. 7 presents this projection for both FIM forecasts and MCMC sampling, showing the full set of viable xSM points as well as those lying within the 68% and 95% confidence regions around the benchmark prediction, thereby enabling a direct comparison between FIM-based and MCMC-based constraints.
Figure 7: Projection of the scanned xSM model points onto the spectral parameter plane $\log_{10}\Omega_{0}$$\log_{10}(f_{p}/\rm Hz)$. The top row corresponds to FIM forecasts, while the bottom row shows results from MCMC sampling. The left panels display scanned parameter points in the geometric parameter space, with blue dots representing the full xSM dataset that survives the phenomenological and theoretical constraints and is capable of generating a first order EWPT. The right panels provide a zoomed-in view of the region surrounding the benchmark prediction. Points falling within the 68% and 95% confidence regions are highlighted in red and green, respectively. This visualization enables a direct comparison between Fisher-based and MCMC-based constraints on GW spectral parameters illustrating how viable model points are distributed with respect to the forecasted confidence contours.
$(v_{s},m_{h_{2}},\theta,b_{3},b_{4})\mapsto(\Omega_{0},f_{\mathrm{p}})$
[…]
Figure 9: Inferred Higgs cubic and quartic couplings $\delta\kappa_{3}$ and $\delta\kappa_{4}$ based on GW measurement of the simulated benchmark signal and MCMC sampling. Blue points represent the full parameter space that survive phenomenological and theoretical constraints and which also can lead to first order EWPT; red and green points highlight subsets that fall within the $1\sigma$ and $2\sigma$ confidence regions, respectively, in the $(\log_{10}\Omega_{0},\,\log_{10}(f_{p}/\rm Hz))$ plane. The black dot denotes the prediction from the fiducial benchmark point. Overlaid horizontal bars indicate the projected $68\%$ and $95\%$ confidence level sensitivities to $\delta\kappa_{3}$ at various future collider experiments, including HL-LHC, CEPC, ILC, FCC, and CLIC taken from Di Vita et al. (2018).
5.4 Measurements of Higgs Self-Couplings
[…]
For all points, due to the features of this model, as explained in more detail in Alves et al. (2018), there appears to be a linear correlation between these two couplings, which can be more easily understood from a Taylor expansion in powers of the mixing angle $\theta$, taking into account the concentrated regions of the parameter space resulting from all phenomenological and theoretical constraints. Comparing the blue points with the green and red points, we can see that the region on the plane $(\delta\kappa_{3},\delta\kappa_{4})$ has narrowed down to a much smaller area due to the GW measurement. This highlights the importance of GW measurement in probing the Higgs couplings, especially regarding the quartic ones. We note, however, that due to the same parameter degeneracy problem, the precision on $\delta\kappa_{3}$ and $\delta\kappa_{4}$ is limited. This limitation, at least for this model, could potentially affect a much broader class of models. A potential solution to this conundrum is to make use of finer structures on the spectrum, such as the damping feature at higher frequencies caused by dissipative effects in the fluid Guo (2023). We further note that in deriving the results here, we have neglected theoretical uncertainties arising from spectrum and phase transition parameter calculations, as explored recently Lewicki et al. (2024). The inclusion of these would lead to a less constraining result. Systematically quantifying and including these uncertainties in the Bayesian framework, along with a more faithful detector simulation, would be a direction for future studies.
6 Conclusion
In this work, we have developed a comprehensive framework that connects the theoretical modeling of FOPTs in the xSM with realistic data analysis strategies for space-based GW detectors, such as Taiji and LISA. By combining frequency-domain detector response modeling, astrophysical foregrounds, and instrumental noise, we constructed simulated datasets for the SGWB and carried out both FIM and Bayesian MCMC analyzes. Using simulated data, we first estimated instrumental and astrophysical parameters, and then extended the analysis to include the stochastic signal from SWs generated during a first-order EWPT. The inferred spectral parameters $(\Omega_{0},f_{\mathrm{p}})$ were subsequently mapped onto the xSM parameter space $(v_{s},m_{h_{2}},\theta,b_{3},b_{4})$, enabling us to identify viable regions consistent with a potential SGWB detection. This mapping further allowed us to translate GW constraints into predictions for Higgs self-coupling deviations $(\delta\kappa_{3},\delta\kappa_{4})$, which encode the shape of the electroweak scalar potential. Our analysis shows that space-based GW observations can help determine these couplings. While collider measurements of $\delta\kappa_{3}$ and especially $\delta\kappa_{4}$ remain extremely challenging, GW-based inference from FOPTs offers an indirect but powerful probe of scalar self-interactions. The synergy between collider physics and GW astronomy thus opens a new pathway toward a more complete understanding of electroweak symmetry breaking and the origin of the Higgs potential. Future work will extend this framework to include detector networks, realistic data gaps, and a joint analysis within the global-fit approach that simultaneously accounts for deterministic sources such as massive black-hole binaries. The inclusion of theoretical uncertainties in the phase-transition modeling—such as gauge dependence and finite-temperature corrections—will also be an important next step toward robust, data-driven constraints on beyond-Standard-Model physics from the upcoming era of space-based GW observations.
Appendix A Fisher Information Matrix and Bayesian Inference
In this appendix, we provide a detailed introduction to the FIM and its formal derivation, followed by a description of the Bayesian inference framework adopted in this work. To illustrate and compare these statistical approaches in a transparent setting, we further present a simple linear toy model as an explicit example of the parameter–inference procedure.
A.1 Fisher Information Matrix and the Cramér–Rao Lower Bound
@@
In this section, we derive the general expression for the FIM. Before doing so, it is necessary to examine the analytical form of the log-posterior functionGregory (2005), which forms the basis for the subsequent derivation:
$\ln\left[p(\bm{\theta})\mathcal{L}(\bm{\theta})\right]=\ln\left[p(\hat{\bm{\theta}})\mathcal{L}(\hat{\bm{\theta}})\right]-\frac{1}{2}\sum_{ij}\left(\theta_{i}-\hat{\theta}_{i}\right)(C^{-1})_{ij}\left(\theta_{j}-\hat{\theta}_{j}\right),$ (A.1)
@@
$F_{ij}\equiv(C^{-1})_{ij}=-E\left(\frac{\partial^{2}\ln p(\bm{\theta})\mathcal{L}(\bm{\theta})}{\partial\theta_{i}\partial\theta_{j}}\right),$ (A.2)
@@
$F_{ij}=F_{ij}^{\text{prior}}+F_{ij}^{\text{likelihood}},$ (A.3)
@@
$\begin{split}F_{ij}^{\text{prior}}&=-E\left(\frac{\partial^{2}\ln p(\bm{\theta})}{\partial\theta_{i}\partial\theta_{j}}\right),\qquad\\ F_{ij}^{\text{likelihood}}&=-E\left(\frac{\partial^{2}\ln\mathcal{L}(\bm{\theta})}{\partial\theta_{i}\partial\theta_{j}}\right).\end{split}$ (A.4)
For an unbiased estimator $\hat{\theta}_{i}$ of an unknown parameter, the CRLB imposes the following lower bound on its standard deviation 94:
$\Delta\hat{\theta}_{i}\geq\sqrt{(F^{-1})_{ii}}.$ (A.5)
@@
To visualize these correlations, one may construct confidence ellipses, which provide a geometric interpretation of the joint uncertainty between two parameters. These ellipses are derived from the second term in Eq. (A.1). The construction begins by analyzing the so-called Mahalanobis distance Mahalanobis (2018):
$d_{M}^{2}=\sum\limits_{ij}\left(\theta_{i}-\hat{\theta}_{i}\right)(C^{-1})_{ij}\left(\theta_{j}-\hat{\theta}_{j}\right),$ (A.6)
@@
$\Sigma=\begin{pmatrix}C_{\alpha\alpha}&C_{\alpha\beta}\\ C_{\alpha\beta}&C_{\beta\beta}\end{pmatrix}$, (A.7)
@@
$\begin{pmatrix}z_{\alpha}\\ z_{\beta}\end{pmatrix}=\Sigma^{-\frac{1}{2}}\begin{pmatrix}\theta_{\alpha}-\hat{\theta}_{\alpha}\\ \theta_{\beta}-\hat{\theta}_{\beta}\end{pmatrix}$, (A.8)
@@
A.2 Bayesian Inference and MCMC Sampling